If the usual adult dose of a drug is 120 mg, what would be the dose based on BSA for a person measuring 6 feet tall and weight 200 lb? (2.2lb/kg; 2.54 cm/in)
Answer #1

Normal Adult dose :120mg

Dose for a person measuring

Height=6 feet=72inch=182.8cm(one feet=12 inch)

Weight=200lb=90.9kg

BSA(body surface area)

m^2=√ht(cm)*wt(kg)/3600

      =√182.8*90.9/3600

      =√4.61

     =2.14meter square

Dosage required=2.14*120/1.7m square

              =151mg

Answer The dose for the person of height 6 feet and 200 pounds body weight is 151mg

Note:√=square root

         ^=raised to